OpenStudy (dumbsearch2):

What is the molecular formula of the compound whose molar mass is 120.2 g/mol and empirical formula is SiO2? @JoannaBlackwelder @aaronq

OpenStudy (dumbsearch2):

This one confuses me even more: A compound is found to be 85.6% C and 14.4% H. Its molecular mass is 42.1 g. What is its molecular formula? A. CH B. CH6 C. C3H6 D. C6H12 Is it C?

OpenStudy (danjs):

For the first one, Silicon atomic mass - about 28.02 g/mol Oxygen atomic mass - about 16 g/mol The empirical formula shows a ratio of one silicon to 2 oxygens, and has a mass of SiO2 - 28.02 + 2(16) = 60.02 g/mol The compound you are looking for has atomic mass 120.2 g/mol \[\frac{ 120.2 }{ 60.02 } \approx 2\] 2x [SiO2] ---> Si2O4

OpenStudy (danjs):

The second one with the percents i am guessing goes something like this ... Carbon - 85.6% of the 42.1g molecular mass 42.1*0.856 = 36.037 g of Carbon molecular mass Molecular mass of carbon is about 12.01 grams per mol \[\frac{ 36.037 }{ 12.01 } = 3~~carbon ~atoms\] Hydrogen - 14.4% of the 42.1 g molecular mass 42.1*0.144 = 6.06 g of Hydrogen Molecular mass Molecular mass of hydrogen is about 1.008 grams per mole \[\frac{ 6.06 }{ 1.008 } \approx 6~hydrogen~atoms\] Overall: \[\large\color{blue}{ C _{3}H _{6}}\]

OpenStudy (dumbsearch2):

@DanJS Does the chemical formula MgCl2 represent a molecule or a formula unit? A. Molecule B. Formula unit